### v1

@@ A.3 Equations for the boundary conditions
$\begin{split}p(z)=&\,~{}p_{0}e^{-i\omega\sqrt{1+b^{2}}z}\left(1-\frac{i(1-ab\omega^{2}+b^{2}\omega^{2})}{\sqrt{1+b^{2}}\omega z}+\mathcal{O}(z^{-2})\right)\\ &+e^{+i\omega bz}\left(\frac{e^{i(a-b)\omega}\sqrt{a-b}\,(\mathcal{R}+\mathcal{T})}{\sqrt{2L}\omega^{2}(1+b^{2})^{1/4}z}+\mathcal{O}(z^{-2})\right)\\ &+e^{-i\omega bz}\left(\frac{e^{-i(a-b)\omega}\sqrt{a-b}}{\sqrt{2L}\omega^{2}(1+b^{2})^{1/4}z}+\mathcal{O}(z^{-2})\right)\,,\end{split}$ (A.9)
@@ Appendix B Briefly review of the transport in conformal interface
$\begin{split}[ds^{2}]_{\text{I}}=&4GL_{\text{I}}\epsilon\left[e^{i\omega(t_{\text{I}}-x_{\text{I}})}(dt_{\text{I}}-dx_{\text{I}})^{2}+\mathcal{R}e^{i\omega(t_{\text{I}}+x_{\text{I}})}(dt_{\text{I}}+dx_{\text{I}})^{2}\right]\,,\\ [ds^{2}]_{\text{II}}=&4GL_{\text{I}}\nu\,\epsilon\mathcal{T}e^{i\omega(t_{\text{II}}-x_{\text{II}})}(dt_{\text{II}}-dx_{\text{II}})^{2}\,,\end{split}$ (B.4)
@@
$\begin{split}t_{\text{I}}=&\,~{}t+\frac{G\epsilon}{L_{\text{I}}}e^{i\omega\tau}\lambda(z)\,,\ x_{\text{I}}=z\tan\alpha+\frac{2G\epsilon}{L_{\text{I}}}e^{i\omega\tau}h_{\text{I}}(z)\,,\ u_{\text{I}}=z+\frac{2G\epsilon}{L_{\text{I}}}e^{i\omega\tau}\xi(z)\,,\\ t_{\text{II}}=&\,~{}t-\frac{G\epsilon}{L_{\text{I}}}e^{i\omega\tau}\lambda(z)\,,\ x_{\text{II}}=-z\nu\tan\beta+\frac{2G\epsilon}{L_{\text{I}}}e^{i\omega\tau}h_{\text{II}}(z)\,,\ u_{\text{II}}=\nu z-\frac{2G\epsilon}{L_{\text{I}}}e^{i\omega\tau}\xi(z)\,,\end{split}$ (B.5)
@@
For the case $L_{\text{I}}=L_{\text{II}}$, the transmission coefficient $\mathcal{T}$ lies in the interval $[\frac{1}{2},1]$. The brane configuration corresponding to $\mathcal{T}=1$ is given by $x_{\text{I}}=x_{\text{II}}=0$, while the brane configuration for $\mathcal{T}=\frac{1}{2}$ coincides with the AdS boundary at $z=0$.

### v2

@@ A.3 Equations for the boundary conditions
[…]
@@ Appendix B Briefly review of the transport in conformal interface
[…]
@@
[…]
@@
For the case $L_{\text{I}}=L_{\text{II}}$, the transmission coefficient $\mathcal{T}$ lies in the interval $\frac{1}{2},1]$. The brane configuration corresponding to $\mathcal{T}=1$ is given by $x_{\text{I}}=x_{\text{II}}=0$, while the brane configuration for $\mathcal{T}=\frac{1}{2}$ coincides with the AdS boundary at $z=0$.
Appendix C Another solution with UV and IR asymptotically AdS2
We consider another brane profile
$\psi(z)=\frac{2a}{b}(\sqrt{1+bz}-1)\,.$ (C.1)
The null energy condition requires $ab>0$ and we again assume $a>0$ and $b>0$. The background brane geometry is asymptotically AdS with radius $\sqrt{1+a^{2}}L$ in the UV and $L$ in the IR. From the UV expansion, it exactly matches (3.1) by identifying that A=a,B=ab/4A=a,B=-ab/4.
The profile for the scalar field is
$\phi=\int_{0}^{z}ds\,\sqrt{\frac{abL}{2s(1+bs)\sqrt{1+a^{2}+bs}}}\,$, (C.2)
and the potential is
$V(\phi(z))=\frac{a(4+4a^{2}+5bz)}{2L(1+a^{2}+bz)^{3/2}}\,.$ (C.3)
The solution to the homogeneous part of ODE (2.22) takes the form
[…]
From the solution (C.4) with $c_{+}=0$, we obtain the transmission coefficient
$\mathcal{T}=1+\frac{4iab\omega}{C_{2}}U\left(\frac{3}{4}+\frac{i(2a^{2}-1)\omega}{2\sqrt{6}b},~\frac{3}{2},~\frac{i\sqrt{6}\omega}{b}\right)\,$, (C.5)
where
[…]
and
$\mathcal{R}=1-\mathcal{T}\,,\ \ c_{-}=-\frac{8iab}{\omega C_{2}}\,.$ (C.7)